OpenStudy (anonymous):

Solve the following problem by working backward. Show and explain each step. Maria received a gift of cash from her grandmother. She put two-thirds of the money in her savings account, spent twenty-five dollars on a pair of jeans, and spent half the remaining money on a DVD. If Maria had eighteen dollars left over, how much money did she receive from her grandmother?

OpenStudy (kamibug):

So working backwards...Cool, I love that! xD Okay, so it wants you to see what the amount of money her grandma gave her. To do that You take the information BACKWARDS! Yaya! So it says at the end that she has $18 left. And before that, it says that she had spent half of what was left after using $25. So, that means you do 18 x 2 = 36. And then do 36 + 25 = 61 since she had spent that before. This means her grandma gave her $61. To check if we're right, you can take $61 and see what she did with it. You need to make sure hse ends up with $18 left. 1. She spent $25. (61-25 = 36.) 2. She spent half of what was left. (36 divided by 2 = 18) 3. 18=18 Yay! =D So, this shows that her grandma gave her $61!!!

OpenStudy (anonymous):

thank you thank you thank you so much @KamiBug...(i love you just joking but thank you) :):):):):):):):):):):):):):):):):):):):):):):):):):):)

OpenStudy (anonymous):

@KamiBug can you help me with this to Write a word problem that can be solved by working backward. Include at least two steps in the problem. Then, show the answer and how the problem can be solved by working backward

OpenStudy (kamibug):

Okay it's definitely gunna take me a while to come up with a problem, lol, hold on...

OpenStudy (anonymous):

okay

OpenStudy (kamibug):

Jesse had a stack of baseball cards. She gave half of her cards to her brother. Then she gave 3 to Penny. Penny said she only wanted one and gave her back 2. Jesse now has 18 cards. How many cards did she have at first?

OpenStudy (anonymous):

@KamiBug what are the steps on doing this

OpenStudy (kamibug):

The equation would be x/2 - 1 = 18. Then you have to solve for x to find the answer to the question. But, that wouldn't be working backwards, though. That would just be making an equation. What do you think the first step is in solving it backwards?